# Estimating Reaction Barriers with Deep Reinforcement Learning

### v1

@@ Abstract
Stable states in complex systems correspond to local minima on the associated potential energy surface. Transitions between these local minima govern the dynamics of such systems. Precisely determining the transition pathways in complex and high-dimensional systems is challenging because these transitions are rare events, and isolating the relevant species in experiments is difficult. Most of the time, the system remains near a local minimum, with rare, large fluctuations leading to transitions between minima. The probability of such transitions decreases exponentially with the height of the energy barrier, making the system’s dynamics highly sensitive to the calculated energy barriers. This work aims to formulate the problem of finding the minimum energy barrier between two stable states in the system’s state space as a cost-minimization problem. We propose solving this problem using reinforcement learning algorithms. The exploratory nature of reinforcement learning agents enables efficient sampling and determination of the minimum energy barrier for transitions.
@@ 1 Introduction
There are multiple sequential decision-making processes one comes across in the world, such as control of robots, autonomous driving, and so on. Instead of constructing an algorithm from the bottom up for an agent to solve these tasks, it would be much easier if we could specify the environment and the state in which the task is considered solved, and let the agent learn a policy that solves the task. Reinforcement learning attempts to address this problem. It is a hands-off approach that provides a feature vector representing the environment and a reward for the actions the agent takes. The objective of the agent is to learn the sequence of steps that maximizes the sum of returns.
@@
One widespread example of a sequential decision-making process where reinforcement learning is utilized is solving mazes Mannucci and van Kampen [2016]. The agent, a maze runner, selects a sequence of actions that might have long-term consequences. Since the consequences of immediate actions might be delayed, the agent must evaluate the actions it chooses and learn to select actions that solve the maze. Particularly, in the case of mazes, it might be relevant to sacrifice immediate rewards for possibly larger rewards in the long term. This is the exploitation-exploration trade-off, where the agent has to learn to choose between leveraging its current knowledge to maximize its current gains or further increasing its knowledge for some possibly larger reward in the long term, possibly at the expense of short-term rewards. The process of learning by an agent while solving a maze is illustrated in Figure 1.
@@
Figure 1: Maze solving using reinforcement learning: (a) The agent is at a state at a particular time step. (b) The agent takes an action (according to the policy it is supposed to learn) and reaches the next state. (c) The agent records the reward obtained by taking the action at that state. (d) The agent takes another action to explore the environment. In most reinforcement learning algorithms this (state, action, reward, next state, next action) tuple is stored while the agent is learning. With a large number of interactions with the environment (e), where the agent takes an action at a state to learn the reward obtained, the agent learns a policy (f) which maximizes the rewards collected by the agent. The policy (f) gives the sequence of actions that the agent has to take from the initial state to the final state so that it collects the maximum rewards in the episode.
@@
GridWorld is an environment for reinforcement learning that mimics a maze Sutton and Barto [1998]. The agent is placed at the start position in a maze with blocked cells, and the agent tries to reach a stop position with the minimum number of steps possible. One might note an analogy of a maze runner with an agent negotiating the potential energy landscape of a transition event for a system along the saddle point with the minimum height. The start state and the stop state are energy minima on the potential energy surface, separated by an energy barrier for the transition. The agent would have to perform a series of perturbations to the system to take it from one minimum (the start state) to another (the end state) through the located saddle point. As in the maze-solving problem, the agent tries to identify the pathway with the minimum energy barrier. If the number of steps is considered the cost incurred in a normal maze, it is the energy along the pathway that is the cost for the transition event. A comparison is attempted in Figure 2.
@@
Figure 2: Estimating reaction barriers by modeling the potential energy surface as a maze: (a) The pathway with the lowest energy barrier as determined by a growing string method on the potential energy surface with 9 intermediate images. (b) The reaction profile, plotted as a solid blue line (interpolated to give a smooth curve) from the pathway determined by the growing string method. The reaction barrier is marked as $\Delta E^{\ddagger}$. (c) The same potential energy surface as in (a) is course grained to construct the maze. The initial state is marked as the yellow box while the final state is marked as the green box. The agent can move to the blue boxes while the red boxes represent walls. The energy cutoff 00 is chosen to determine whether a grid box is a wall or not. (d) Instead of the extreme classification of a grid box as a wall or not, each grid box is assigned an energy value. The agent is then allowed to learn the path with the minimum sum of energies along it’s path from the initial state to the final state. The energy profile for the pathway identified by the agent in this maze is plotted as the dashed green line in (b). As it might be seen, coarse-graining the potential energy surface into a $8\times 8$ maze, ans solving then solving it using standard reinforcement learning algorithms provides a reasonable starting point to solving the problem.
@@
Noting the superficial similarities between solving a maze and determining the transition pathway with the lowest energy barrier, we propose to use standard and tested deep reinforcement learning algorithms used to solve mazes in an attempt to solve the problem of finding minimum energy pathways. The problem is formulated as a min-cost optimization problem in the state space of the system. We use this formulation to determine the barrier height of the optimal pathway in the Mueller-Brown potential. Neural nets are used as the actor and critic function approximators, and a randomly perturbed policy is used to facilitate exploration of the potential energy surface by the agent. Delayed policy updates and target policy averaging are used to stabilize the learning, especially during the first few epochs, which are crucial to the optimal performance of the agent.
Section 2 describes the methods used to formulate the problem as a Markov decision process and the algorithm used to solve it. Section 3 elaborates on the experiments where the formulated method is used to determine the barrier height of a transition on the Mueller-Brown potential. Section 4 contains a short discussion of the work in the context of other similar studies and the conclusions drawn from this work.
@@ 2 Methods
To solve the problem of finding a pathway with the lowest energy barrier for a transition using reinforcement learning, one has to model it as a Markov decision process. Any Markov decision process consists of (state, action, next state) tuples. In this case, the agent starts at the initial state (a local minimum) and perturbs the system (the action) to reach a new state. Since the initial state was an energy minimum, the current state will have higher energy. However, as in many sequential control problems, the reward is delayed. A series of perturbations that lead to states with higher energies might enable the agent to climb out of the local minimum into another one containing the final state. By defining a suitable reward function and allowing the agent to explore the potential energy surface, it is expected that the agent will learn a path from the initial to the final state that maximizes the rewards. If the reward function is defined properly, it should correspond to the pathway with the lowest energy barrier for the transition.
Once the problem is formulated as a Markov decision process, it can be solved by some reinforcement learning algorithm. Twin Delayed Deep Deterministic Policy Gradient (TD3) Fujimoto et al. [2018] is a good start because it prevents the overestimation of the state value function, which often leads to the agent exploiting the errors in the value function and learning a sub-optimal policy. Soft Actor Critic (SAC) Haarnoja et al. [2018] tries to blend the deterministic policy gradient with a stochastic policy optimization, promoting exploration by the agent. In practice, using a stochastic policy to tune exploration often accelerates the agent’s learning.
@@ 2.1 Markov Decision Process
At a state $s^{(k)}$, the agent takes an action $\Delta s^{(k)}$. Since the action is considered a perturbation to the current state of the system, the next state $s^{(k+1)}$ is determined from the current state $s^{(k)}$ as $s^{(k+1)}=s^{(k)}+\lambda\cdot\Delta s^{(k)}$.
@@
Since both the state space and action space are continuous, an actor-critic based method, specifically the soft actor-critic (SAC), is used. Additionally, since the state space is continuous, the episode is deemed to have terminated when the difference between the current state and the target state is smaller than some tolerance, $x\in\mathbb{R}^{d}:|x-x_{t}|<\delta$ for some small $\delta$. Otherwise, it would be extremely unlikely that the agent would land exactly at the coordinates of the final state after taking some action.
@@ 2.2 Algorithm
4:training_step = 1
5:for step = 1 to numEpisodes do
6:     state, _ = env.reset()
7:     for t = 0 to maxSteps do
8:         let actor select an action by policy $\pi_{\theta}$
9:         perturb the action with some noise $\epsilon\sim\mathcal{N}(0,\sigma)$
$a^{(k)}=\pi_{\theta}(s^{(k)})+\text{clip}(\epsilon,-\epsilon_{\text{lim}},\epsilon_{\text{lim}})$
10:         execute action in the env and observe the $\left(s^{(k)},a^{(k)},r^{(k)},s^{(t+1)},\text{{ done}}\right)$ tuple
11:         push it to the replay buffer $\mathcal{R}$
12:         if training_step % agent_update == 0  then
13:              sample a minibatch $\mathcal{B}$ of $\left(s,a,r,s^{\prime},\text{{ done}}\right)$ tuples from the replay buffer $\mathcal{R}$
14:              compute targets as (where $a^{\prime}=\pi_{\theta}(\cdot|s^{\prime})+\text{clip}(\epsilon,-\epsilon_{\text{lim}},\epsilon_{\text{lim}})$)
$y(r,s^{\prime})=r+\gamma(1-\text{{done}})\left(\min_{i=1,2}Q_{\phi_{i,\text{target}}}(s^{\prime},a^{\prime})-\alpha\log\pi_{\theta}(a^{\prime}|s^{\prime}))\right)$
15:              update critic Q-nets parameters by one step of gradient descent with loss function (with the gradient clipped by some maximum value)
@@
16:              if t % update_target == 0 then
17:                  update actor net parameters by one step of gradient descent with loss function (with the gradient clipped by some maximum value)
$\frac{1}{|\mathcal{B}|}\,\,\,\nabla_{\theta}\left(\sum_{s\in\mathcal{B}}\min_{i=1,2}Q_{\phi_{i}}(s,\pi_{\theta}(s))-\alpha\log\pi_{\theta}(a|s)\right)^{2}$
18:                  update the entropy coefficient $\alpha$ as one step of gradient descent with loss function
$\frac{1}{|\mathcal{B}|}\,\,\,\nabla_{\alpha}\left(\sum_{s\in\mathcal{B}}-\alpha\log\pi_{\theta}(a|s)-\alpha\log a^{\prime}\right)^{2}$
19:                  soft update the target networks: $\phi_{i,\text{target}}\leftarrow\tau\phi_{i}+(1-\tau)\phi_{i,\text{target}}$
@@
23:end for
24:return the actor net parameters $\theta$ and critic Q-net parameters $\phi_{1},\phi_{2}$.
SAC, an off-policy learning algorithm with entropy regularization, is used to solve the formulated Markov Decision process because the inherent stochasticity in its policy facilitates exploration by the agent. The algorithm learns a behavior policy $\pi_{\theta}$ and two critic Q-functions, which are neural nets with parameters $\phi_{1}$ and $\phi_{2}$ (line 1 of Algorithm 1).
The agent chooses an action $a^{(k)}\equiv\Delta s^{(k)}$ to take when at state $s^{(k)}$ following the policy $\pi_{\theta}$ (line 8). The returns from the state $s^{(k)}$ when acting according to the policy $\pi$ is the discounted sum of rewards collected from that step onwards till the end of the episode: $R_{t}=-\sum_{i=t}^{T}\gamma^{i-t},E(s^{(i)})$. The objective of the reinforcement learning agent is to determine the policy $\pi^{*}$ that maximizes the returns, $R_{t}$, for states $s\in\mathcal{S}$. This is done by defining a state-action value function, $Q(s^{(i)},a^{(i)})$, which gives an estimate of the expected returns if action $a^{(i)}$ is taken by the agent when at state $s^{(i)}$: $Q(s^{(i)},a^{(i)})=\mathbb{E}\left[R_{t}:s_{t}=s^{(i)},a_{t}=a^{(i)}\right]$. Since the objective is to maximize the sum of the returns, the action-value function can be recursively defined as
$Q(s^{(i)},a^{(i)})=-E(s^{(i+1)})+\max_{a^{(i+1)}\in\mathcal{A}}Q(s^{(i+1)},a^{(i+1)})$
@@
The entropy coefficient $\alpha$ is adjusted over the course of training to encourage the agent to explore more when required and exploit its knowledge at other times (line 18) Haarnoja et al. [2019]. However, we also borrow some elements from the TD3 algorithm Fujimoto et al. [2018] to improve the learning of the agent, namely delayed policy updates and target policy smoothing. The critic Q-nets are updated more frequently than the actor and the target Q-nets to allow the critic to learn faster and provide more precise estimates of the returns from the current state. To address the problem of instability in the learning, especially in the first few episodes while training the agent, target critic nets are used. Initially, the critic nets are duplicated (line 2) and subsequently soft updates of these target nets are carried out after an interval of a certain number of steps (line 19). This provides more precise estimates for the state-action value function while computing the returns for a particular state in line 14. To encourage the agent to explore the potential energy surface, clipped noise is added to the action chosen by the actor net (line 9). This also makes it difficult for the actor to exploit imprecise Q-net estimates during the beginning of the training. The changes to the SAC algorithm, borrowed from TD3, are highlighted in blue in the pseudocode of Algorithm 1. The parameters used in the particular implementation of the algorithm are listed in Table 1.
@@
scaling factor for actions $0.015$
@@
maximum steps per epoch $500$
number of training epochs $1000$

### v2

@@ Abstract
Stable states in complex systems correspond to local minima on the associated potential energy surface. Transitions between these local minima govern the dynamics of such systems. Precisely determining the transition pathways in complex and high-dimensional systems is challenging because these transitions are rare events, and isolating the relevant species in experiments is difficult. Most of the time, the system remains near a local minimum, with rare, large fluctuations leading to transitions between minima. The probability of such transitions decreases exponentially with the height of the energy barrier, making the system’s dynamics highly sensitive to the calculated energy barriers. This work aims to formulate the problem of finding the minimum energy barrier between two stable states in the system’s state space as a cost-minimization problem. It is proposed to solve this problem using reinforcement learning algorithms. The exploratory nature of reinforcement learning agents enables efficient sampling and determination of the minimum energy barrier for transitions.
@@ 1 Introduction
There are multiple sequential decision-making processes that one comes across in the world, such as control of robots, autonomous driving, and so on. Instead of constructing an algorithm from the bottom up for an agent to solve these tasks, it would be much easier if one could specify the environment and the state in which the task is considered solved and let the agent learn a policy that solves the task Kaelbling et al. [1996], Maes [1993]. Reinforcement learning attempts to address this problem. It is a hands-off approach that provides a feature vector representing the environment and a reward for the actions the agent takes Sutton and Barto [1998]. The objective of the agent is to learn the sequence of steps that maximizes the sum of returns. Silver et al. [2021]
@@
One widespread example of a sequential decision-making process in which reinforcement learning is utilized is solving mazes Mannucci and van Kampen [2016], Osmanković and Konjicija [2011]. The agent, a maze runner, selects a sequence of actions that might have long-term consequences Wiering and van Hasselt [2008]. Since the consequences of immediate actions might be delayed, the agent must evaluate the actions it chooses and learn to select actions that solve the maze. Particularly, in the case of mazes, it might be relevant to sacrifice immediate rewards for possibly larger rewards in the long term. This is the exploitation-exploration trade-off, where the agent has to learn to choose between leveraging its current knowledge to maximize its current gains or further increasing its knowledge for a potential larger reward in the long term, possibly at the expense of short-term rewards Brunner et al. [2018], Parisotto and Salakhutdinov [2017]. The process of learning by an agent while solving a maze is illustrated in Figure 1.
@@
Figure 1: Maze solving using reinforcement learning: (a) The agent is at a state at a particular time step, and takes an action to reaches the next state (b). The agent records the reward obtained by taking the action in that state and (c) continues exploring the environment. After a large number of interactions with the environment, the agent learns a policy (d) that maximizes the rewards collected by the agent. The policy (d) gives the sequence of actions that the agent has to take from the initial state to the final state so that it collects the maximum rewards in an episode.
@@
GridWorld is an environment for reinforcement learning that mimics a maze Sutton and Barto [1998]. The agent is placed at the start position in a maze with blocked cells, and the agent tries to reach a stop position with the minimum number of steps possible. One might note an analogy of a maze runner with an agent negotiating the potential energy landscape of a transition event for a system along the saddle point with the minimum height. The start state and the stop state are energy minima on the potential energy surface, separated by an energy barrier for the transition. The agent would have to perform a series of perturbations to the system to take it from one minimum (the start state) to another (the end state) through the located saddle point. In the case of a maze, the (often discrete) action changes the position of the agent (by a fixed measure), but while locating minimum energy pathways, the physical problem demands a continuous action space. However, as in the case of a maze, an action changes the variables describing the system (be it physical coordinates or state variables) by a small measure.
As in the maze-solving problem, the agent tries to identify the pathway with the minimum energy barrier. If the number of steps is considered to be the cost incurred in a normal maze, then it is the energy along the pathway that is the cost of the transition event. Reward maps can vary depending on the maze considered, but in the original GridWorld problem, the agent was given a negative reward if the action led to a wall cell, and a zero reward for all non-terminal states (a discount factor $<1$ enforces the minimum count of steps). In the case of locating trajectories with a low energy barrier, the agent should be penalized if the action leads to a state with progressively higher energies (but not to an extent that it hinders exploration). The exact reward map used is detailed later in Section 2.
A comparison is attempted in Figure 2. A smooth potential energy surface is coarse-grained to construct a maze, where all positions with a negative potential energy are shaded blue (possible move cells), while those with a positive potential energy are shaded red (representing walls). The initial state in the maze is marked yellow, while the final state is marked green. However, instead of classifying a grid cell of the constructed maze either as a wall or as a cell, one can discretize the state space and assign an energy value to a cell. An agent can then be trained to reach the final state starting from the initial state and collect the maximum sum of rewards along its path (minimizing the energy along the pathways requires assigning the negative of the energy as the reward for an action leading the agent to the cell). Since this is an episodic problem, one already runs into the problem where the agent moves back and forth between two adjacent cells, collecting rewards from each move in an attempt to maximize the sum of rewards collected, rather than reaching the final state and terminating the episode. For this simple setting, the problem is solved by rewarding the agent only the first time it visits a cell and terminating the episode after a fixed number of steps (in this case, $15$). The energy profile of the pathway followed by the agent (inferred from the rewards collected in an episode) in this maze is plotted as the dashed green line in Figure 2b. As can be seen, coarse-graining the potential energy surface into an $8\times 8$ maze and then solving it using standard reinforcement learning algorithms provides a reasonable starting point to address the problem.
@@
Figure 2: Estimating reaction barriers by modeling the potential energy surface as a maze: (a) The pathway with the lowest energy barrier as determined by a growing string method on the potential energy surface with 9 intermediate images. (b) The reaction profile, plotted as a solid blue line (interpolated to give a smooth curve) from the pathway determined by the growing string method. The reaction barrier is marked as $\Delta E^{\ddagger}$. Instead of the extreme binary classification of a grid cell as a wall or move as in the maze (c), each cell can be assigned an energy value as in (d).
@@
Noting the superficial similarities between solving a maze and determining the transition pathway with the lowest energy barrier, it is proposed to use standard and tested deep reinforcement learning algorithms used to solve mazes in an attempt to solve the problem of finding minimum energy pathways. The problem is formulated as a min-cost optimization problem in the state space of the system. An actor function approximator suggests the action to be taken by the agent when it is in a particular state. A critic function approximator provides an estimate of the sum of rewards until the end of the episode from the new state after taking the action suggested by the actor. Actor-critic based reinforcement learning techniques have been shown to solve problems effectively, even in higher dimensions Zhang [2020]. Neural nets are used as the actor and critic function approximators, and a randomly perturbed policy is used to facilitate exploration of the potential energy surface by the agent. Delayed policy updates and target policy averaging are used to stabilize learning, especially during the first few epochs, which are crucial to the optimal performance of the agent. This formulation is used to determine the barrier height of the optimal pathway in the Müller-Brown potential.
Section 2 describes the methods used to formulate the problem as a Markov decision process and the algorithm used to solve it. Section 3 elaborates on the experiments in which the formulated method is used to determine the barrier height of a transition on the Müller-Brown potential. Section 4 contains a short discussion of the work in the context of other similar studies, while Section 5 outlines the conclusions drawn from this work.
@@ 2 Methods
To solve the problem of finding a pathway with the lowest energy barrier for a transition using reinforcement learning, one has to model it as a Markov decision process. Any Markov decision process consists of (state, action, next state) tuples. In this case, the agent starts at the initial state (a local minimum) and perturbs the system (action) to reach a new state. Since the initial state was an energy minimum, the current state will have a higher energy. However, as in many sequential control problems, the reward is delayed. A series of perturbations that lead to states with higher energies might enable the agent to climb out of the local minimum into another one that contains the final state. By defining a suitable reward function and allowing the agent to explore the potential energy surface, it is expected that the agent will learn a path from the initial state to the final state that maximizes the rewards. If the reward function is defined properly, it should correspond to the pathway with the lowest energy barrier for the transition.
Once the problem is formulated as a Markov decision process, it can be solved by some reinforcement learning algorithm. In most reinforcement learning algorithms, this (state, action, reward, next state, next action) tuple is stored while the agent is learning. Twin Delayed Deep Deterministic Policy Gradient (TD3) Fujimoto et al. [2018] is a good start because it prevents overestimation of the state value function, which often leads the agent to exploit the errors in the value function and learn a suboptimal policy. Soft Actor Critic (SAC) Haarnoja et al. [2018] tries to blend the deterministic policy gradient with a stochastic policy optimization, promoting exploration by the agent. In practice, using a stochastic policy to tune exploration often accelerates the agent’s learning.
@@ 2.1 Markov Decision Process
In a state $s^{(k)}$, the agent takes an action $\Delta s^{(k)}$. Since the action is considered a perturbation of the current state of the system, the next state $s^{(k+1)}$ is determined from the current state $s^{(k)}$ as $s^{(k+1)}=s^{(k)}+\lambda\cdot\Delta s^{(k)}$.
@@
Since both the state space and action space are continuous, an actor-critic based method, specifically the soft actor-critic (SAC), is used. Additionally, since the state space is continuous, the episode is deemed to have terminated when the difference between the current state and the target state is smaller than some tolerance, $x\in\mathbb{R}^{d}:|x-x_{t}|<\delta$ for some small $\delta$. Otherwise, it would be extremely unlikely that the agent would land exactly at the coordinates of the final state after taking some action. An obvious problem with this definition of the Markov process is that the agent may prefer to remain in a state near the target state (but far enough so that the episode does not terminate), collecting rewards for the rest of the episode. This behavior was observed in Section 3.
@@ 2.2 Algorithm
4:for step = 1 to numEpisodes do
5:     state, _ = env.reset()
6:     for t = 0 to maxSteps do
7:         let actor select an action by policy $\pi_{\theta}$
8:         perturb the action with some noise $\epsilon\sim\mathcal{N}(0,\sigma)$ : $a^{(k)}=\pi_{\theta}(s^{(k)})+\text{clip}(\epsilon,-\epsilon_{\text{lim}},\epsilon_{\text{lim}})$
9:         execute action in the env and observe the $\left(s^{(k)},a^{(k)},r^{(k)},s^{(t+1)},\text{{ done}}\right)$ tuple
10:         push it to the replay buffer $\mathcal{R}$
11:         if step % agent_update == 0  then
12:              sample a minibatch $\mathcal{B}$ of $\left(s,a,r,s^{\prime},\text{{ done}}\right)$ tuples from the replay buffer $\mathcal{R}$
13:              compute targets as (where $a^{\prime}=\pi_{\theta}(\cdot|s^{\prime})+\text{clip}(\epsilon,-\epsilon_{\text{lim}},\epsilon_{\text{lim}})$)
 $\text{\hskip 56.9055pt}y(r,s^{\prime})=r+\gamma(1-\text{{done}})\left(\min_{i=1,2}Q_{\phi_{i,\text{target}}}(s^{\prime},a^{\prime})-\alpha\log\pi_{\theta}(a^{\prime}|s^{\prime}))\right)$
14:               update critic Q-nets parameters by one step of gradient descent with loss function (with the gradient clipped by some maximum value)
@@
15:              if t % update_target == 0 then
16:                   update actor net parameters by one step of gradient descent with loss function (with the gradient clipped by some maximum value)
 $\text{\hskip 71.13188pt}\frac{1}{|\mathcal{B}|}\,\,\,\nabla_{\theta}\left(\sum_{s\in\mathcal{B}}\min_{i=1,2}Q_{\phi_{i}}(s,\pi_{\theta}(s))-\alpha\log\pi_{\theta}(a|s)\right)^{2}$
17:                  update the entropy coefficient $\alpha$ as one step of gradient descent with loss function
 $\text{\hskip 71.13188pt}\frac{1}{|\mathcal{B}|}\,\,\,\nabla_{\alpha}\left(\sum_{s\in\mathcal{B}}-\alpha\log\pi_{\theta}(a|s)-\alpha\log a^{\prime}\right)^{2}$
18:                  soft update the target networks: $\phi_{i,\text{target}}\leftarrow\tau\phi_{i}+(1-\tau)\phi_{i,\text{target}}$
19:              end if
@@
23:return the actor net parameters $\theta$ and critic Q-net parameters $\phi_{1},\phi_{2}$.
SAC, an off-policy learning algorithm with entropy regularization, is used to solve the formulated Markov Decision process because the inherent stochasticity in its policy facilitates exploration by the agent. Entropy regularization tries to balance maximizing the returns till the end of the episode with randomness in the policy driving the agent. The algorithm learns a behavior policy $\pi_{\theta}$ and two critic Q-functions, which are neural networks with parameters $\phi_{1}$ and $\phi_{2}$ (line 1 of Algorithm 1).
The agent chooses an action $a^{(k)}\equiv\Delta s^{(k)}$ to take when at state $s^{(k)}$ following the policy $\pi_{\theta}$ (line 8). Returns from state $s^{(k)}$ when acting according to policy $\pi$ is the discounted sum of rewards collected from that step onward to the end of the episode: $R_{t}=-\sum_{i=t}^{T}\gamma^{i-t}\,E(s^{(i)})$. The objective of the reinforcement learning agent is to determine the policy $\pi^{*}$ that maximizes the returns $R_{t}$, for states $s\in\mathcal{S}$. This is done by defining a state-action value function, $Q(s^{(i)},a^{(i)})$, which gives an estimate of the expected returns if the agent takes action $a^{(i)}$ when in state $s^{(i)}$: $Q(s^{(i)},a^{(i)})=\mathbb{E}\left[R_{t}:s_{t}=s^{(i)},a_{t}=a^{(i)}\right]$. Since the objective is to maximize the sum of the returns, the action-value function can be recursively defined as
$Q(s^{(i)},a^{(i)})=-E(s^{(i+1)})+\gamma\,\max_{a^{(i+1)}\in\mathcal{A}}Q(s^{(i+1)},a^{(i+1)})$
@@
The entropy coefficient $\alpha$ is adjusted over the course of training to encourage the agent to explore more when required and to exploit its knowledge at other times (line 18) Haarnoja et al. [2019]. However, some elements from the TD3 algorithm Fujimoto et al. [2018] are borrowed to improve the learning of the agent, namely delayed policy updates and target policy smoothing. Due to the delayed policy updates, the critic Q-nets are updated more frequently than the actor and the target Q-nets to allow the critic to learn faster and provide more precise estimates of the returns from the current state. To address the problem of instability in learning, especially in the first few episodes while training the agent, target critic nets are used. Initially, the critic nets are duplicated (line 2), and subsequently soft updates of these target nets are carried out after an interval of a certain number of steps (line 19). This provides more precise estimates for the state-action value function while computing the returns for a particular state in line 14. Adding noise to input during the training of a machine learning model often leads to improved performance because it acts as an $L_{2}$-regularizer Bishop [1995] and prevents the model from memorizing patterns in the training data sample in supervised learning scenarios. Similarly, TD3 adds noise to the action predicted by the actor network to smooth out the Q-function, so that the agent does not memorize the imprecise estimates of the Q-function early on during the training process. This prevents the policy from exploiting the imprecise estimates of the Q-function approximator for certain actions, reducing the chances of learning a brittle policy that does not generalize well. The logic is that, for well-behaved, smooth reward maps, the reward should not abruptly change with small differences in the action. The addition of clipped noise to the action chosen by the actor net (in line 9) also encourages the agent to explore the potential energy surface. The changes to the SAC algorithm, borrowed from TD3, are highlighted in blue in the pseudocode of Algorithm 1. The parameters used in the particular implementation of the algorithm are listed in Table 1.
@@
$\lambda$ or scaling factor for actions $0.01$
@@
maximum steps per episode $500$
number of training epochs $10000$